OpenStudy (anonymous):

find the solution- cosxcos2xcos3x=1/4

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

emm maybe writing\[\cos x\cos 3x\]formula will help

OpenStudy (anonymous):

\[\cos a \cos b=?\]

OpenStudy (anonymous):

2cosacosb ??

OpenStudy (anonymous):

=cos(a+b)+cos(a-b)

OpenStudy (anonymous):

oops that will not help ??

OpenStudy (anonymous):

lol